# Alignment-Aware Model Extraction Attacks on Large Language Models

### v1

@@ 6 Experiments
    Model/Metric Accuracy Precision Recall F1 Score
PIQA [10] with 64 query samples
Victim Model 0.828 0.828 0.827 0.827
Local Model 0.622 0.638 0.621 0.609
+LoRD (ours) $0.785\pm 0.01$ $0.795\pm 0.01$ $0.785\pm 0.01$ $0.783\pm 0.02$
TruthfulQA [33] with 64 query samples
Victim Model 0.414 0.500 0.207 0.293
Local Model 0.391 0.500 0.195 0.281
+LoRD (ours) $0.408\pm 0.05$ $0.500\pm 0.00$ $0.204\pm 0.03$ $0.289\pm 0.03$
Table 3: MEA comparison on QA tasks among MLE and our LoRD methods, where we use GPT-3.5-turbo as the victim model, and Llama3-8B [1] as the local initial model.
@@ 6.1 Settings
Datasets. We evaluate MEAs on five mainstream natural language generation (NLG) tasks, including machine translation, text summarization, question answering, structured text generation, and data-to-text. We select ten representative datasets: WMT16 [11], TLDR [27], CNN Daily Mail [20], Samsum [17], WikiSQL [58], Spider [53], E2E-NLG [14], CommonGen [31], PIQA [10], and TruthfulQA [33] as benchmarks for our domain-specific evaluation. These datasets cover most of the downstream tasks in natural language generation. We compare not only the stealing efficacy of different MEA methods, but also the stealing difficulty across different downstream tasks. Table 1 lists all datasets and backbones used in the paper.
Baselines. As described in Section 2.2 and 5.1, we compare LoRD with two types of model extraction methods: maximum likelihood estimation (MLE) and knowledge distillation (KD). For MLE and LoRD, we conduct MEAs under pure black-box attack settings. For KD, the predicted distributions are used specifically under grey-box settings.
Metrics. For text generation tasks, we evaluate extracted models with a semantic-level and two lexical-level metrics, BERTScore [55], BLEU [38], and Rouge-L [32], all of which are commonly used in the NLG evaluation. Regarding reasoning tasks (e.g. QA), we use Precision, Recall, Accuracy, and F1 score as their evaluation metrics.
Implementation Details. We use Llama3-8B as the local model to learn the outputs generated by victim models. We set sequence length varying 128 to 4096 depending on the selected tasks, and learning rate $3\times 10^{-5}$. Our experiments run on 2 $\times$ 80GB Nvidia Tesla A100. We execute each training five times and record the mean values and standard variances in the following sections. For LoRD, we set $\tau_{1}$ and $\tau_{2}$ to 0.8 and -0.1, respectively. Besides, we set the period number $N_{t}$ to 512, and use $\lambda_{1}=0.5$ as the default formation of the loss function. The victim model’s response, $\mathbf{y}_{vic}$, is generated by token sampling with a temperature of 1, the default setting for OpenAI APIs. The local model also uses token sampling, but with a temperature of 0.8 and Top-P probability clipping [22] at 0.98. We use this setting to enhance the stability of generation in local models. Note that we have not incorporated sampling strategies with their corresponding hyperparameters into the design of LoRD. We believe that MEAs considering sampling strategies could inspire more powerful MEA methods, and we leave these improvements for future work.
@@
Table 4: MEA comparison on WMT16 [11] among MLE and our LoRD methods, where we use GPT-3.5-turbo as the victim model, and Llama3-8B [1] as the local initial model.
@@ 6.2 Stealing Domain-Specific Knowledge
We first select GPT-3.5-turbo, a checkpoint of ChatGPT, as the basic victim model. This is because its API provides probabilities of candidate words when generating responses. We employ Llama3-8B [1], a small LLM with only a 4.5% fraction of parameters than the victim model as our initial local model. Though this LaViSH setting contradicts previous assumptions [50, 37, 23] in MEA that the copy model should usually be “wider” or “larger” than the victim model to contain its knowledge, we believe this setting is more applicable in real-world scenarios [30]. Besides, the number of query times selected in this section is less than 100, a significant degradation compared to previous studies [30]. This is because, in our experiments, copy models can easily learn the knowledge with a few training samples and then exhibit only slight improvements afterward. More discussions on query times can be found in Section 7.2.1.
@@
Fidelity and limits on stealing. We first examine the fidelity and limits of a small LLM to steal commercial LLMs. As shown in Table 2, we list the performance of the victim model and the local model on three tasks, and provide two MEA methods, local model fine-tuned with MLE (+MLE) and LoRD (+LoRD), respectively. In table 2, cells highlighted in red indicate poorer outcomes compared to the victim model, whereas blue signifies results that are on par or potentially superior to the victim model. The intensity of the red or blue color corresponds to the degree of underperformance or outperformance relative to the victim model.
@@
We can see that the original performance of the local model is significantly lower than the victim model, i.e., with a 50% decrease in BLEU-4 or $10\sim 25$ decrease in Rouge-L. Once we employ MEAs in the local model, its performance rapidly boosts to nearly the same as the victim model, with $0\sim 40\%$ points of gaps in BERTScore. These gaps are negligible (e.g. $<1\%$ in summarization) in some tasks, but remain eminent in other tasks such as reasoning, structured text generation, and machine translation. This phenomenon indicates that domain-specific model extractions can effectively learn domain-specific abilities from victim models but may perform poorly if downstream tasks require extra knowledge, such as machine translation and QA.
Figure 6: Spectrum of the fidelity and performance-up on extracting different downstream tasks. Current datasets can be divided into three groups: high fidelity and high performance-up (HFHP), high fidelity but low performance-up (HFLP), and low fidelity but high performance-up (LFHP).
@@
Comparison among stealing methods. Tables 2, 3, and 4 compare the stealing efficacy between MLE and our LoRD. The results consistently show that LoRD significantly outperforms MLE under the same MEA settings. Besides, for challenging tasks such as reasoning and translation, LoRD exhibits much higher improvements, which demonstrates that it can address the preference overfitting problem discussed in Section 5.2 and do enable the local model to learn the task ability from victim models. However, we also observe that for some tasks (e.g. summarization), LoRD shows no statistical difference from MLE, probably because these tasks are relatively simple, where merely MLE has already achieved comparable results to victim models.
@@
where $\mathbf{y}_{N_{t}}\sim P_{\theta_{N_{t}}}(\cdot|\mathbf{x})$, $\mathbf{y}_{0}\sim P_{\theta_{0}}(\cdot|\mathbf{x})$, and $\mathbf{y}_{vic}\sim P_{\theta_{vic}}(\cdot|\mathbf{x})$ denote the sampled responses from the trained local model ($\theta_{N_{t}}$), the initial local model ($\theta_{0}$), and the victim model ($\theta_{vic}$), respectively. In Figure 6, we illustrate a “spectrum” of extracting various downstream tasks based on these two metrics defined in Equation 18. The figure can assist in recognizing and defending commercial LLM’s knowledge.
From Figure 6, we observe five tasks forming the following three scenario groups and datasets coming from the same tasks are mostly in the same group:
@@
  • High fidelity and high performance-up (HFHP): These tasks are challenging for a pre-trained model but can be effectively learned with the guidance of victim models. This group includes two tasks: data-to-text and structured text generation.
@@
  • High fidelity but low performance-up (HFLP): The initial local model already achieves a comparable performance to the victim model. QAs and summarization are in this group.
  • Low fidelity but high performance-up (LFHP): While MEAs significantly improve the local model’s performance, gaps between the local and victim models remain difficult to bridge with domain-specific extraction alone. Machine translation is a representative task whose reasons are explained in Section 6.2.
Figure 7: Comparison of the resistance of watermarks.
@@
7 Empirical Analysis
7.1 Resistance to Watermarks
Current LLM watermarking methods have been shown [26] to be robust against commonly used erasing strategies (e.g., rephrasing), making watermark removal a distinct challenge. In this section, we validate the inherent resistance of LoRD to watermarks, suggesting that LoRD is preliminarily resistant to text watermarking. As described in Section 5, we highlight that LoRD can extract the victim models’ knowledge with two terms: the straightforward likelihood learning term $\text{log}P_{\theta_{t}}(\mathbf{y}_{vic}|\mathbf{x})-\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})$ and the exploration term $\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|\mathbf{x})-\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})$, where we can tune the hyperparameter $\lambda_{1}$ as shown in Equation 17 to trade off the exploration and the convergence speed. Typically, a lower $\lambda_{1}$ encourages the model for conducting a slower but more diverse and localized exploration from its own generated text $\mathbf{y}_{t-1}^{+}$, potentially enhancing watermark resistance. In this subsection, we evaluate this analysis empirically.
Watermarking Details. Unlike previous experimental settings in Section 6, here we cannot utilize commercial LLMs as victim models due to the inability to control token sampling inside LLMs. Instead, we employ Llama3-70B as the victim model and watermark its outputs based on “green” tokens selection. Following prior research [25], we separate the predicted vocabulary into a green word set and a red word set , assigning them randomly with the seed derived from the hash of generated tokens at the last generation step. Subsequently, we sample the next token exclusively from the green set, determined by a certain probability.
7.2 Scaling Laws of LLMs
We then explore essential capacities to steal domain-specific knowledge from LLMs. We first analyze the influence of query times for the adversary, then compare the efficacy when utilizing different sizes of the local model, and finally compare the fidelity among different victim and local models.
7.2.1 Query Times
@@
7.2.2 Scales of Local Models
As discussed in Section 3, we assume the adversary is stealing existing commercial LLMs with a small local model. This raises the question of selecting an appropriate interval of the local model’s size. To address this concern, we illustrate the correlation between the local model’s size and extraction efficacy on two machine translation tasks, Russian-to-English (ru-en) and German-to-English (de-en), as shown in Figure 9. Here, we employ seven OPT models [54] as local models, with parameters ranging from 125 million to 30 billion, to minimize the interruptions of factors other than model size.
@@ 7.2.2 Scales of Local Models
7.2.3 Fidelity under Different Victim and Local Models
We then evaluate the fidelity of extracting different victim models using various pre-trained local models. Specifically, we select GPT-3.5, GPT-4, and GPT-4o as victim models, and employ five state-of-the-art open-source models, Phi-3 (3.8B), OPT (6.7B), Qwen-2 (7B), Mistral-V3 (7B), and Llama-3 (8B), as local models, as shown in Figure 10.
Horizontally, while GPT-4 exhibits a consistently lower extracted fidelity compared to the other two victim models, vulnerabilities of the three victim models are generally similar. Vertically, fidelity of different local models can be significantly impacted by their performance. For instance, OPT (6.7B) shows a noticeably lower score compared to the other four models, which indicates that the initial performance of the local model will affect the performance of MEAs. Besides, Phi-3 (3.8B) achieves a comparable fidelity to larger models like Llama-3 (8B), demonstrating that the size of a local model does not influence final fidelity in domain-specific stealing after 2.7 billion, which corroborates the observation in Section 7.2.2.

### v2

@@ 5.1 Settings
Datasets. We evaluate MEAs on six mainstream natural language generation (NLG) tasks, including safety alignment, machine translation, text summarization, question answering, structured text generation, and data-to-text. We select twelve representative datasets, including two for safety alignment and ten for domain-specific evaluation, as detailed in Table 9. We believe these datasets encompass the majority of downstream tasks and effectively capture the varying degrees of difficulty in model stealing across different task domains.
Baselines. As described in Section 2.2 and 4.1, we compare LoRD with two types of model extraction methods: maximum likelihood estimation (MLE) and knowledge distillation (KD). For MLE and LoRD, we conduct MEAs under pure black-box attack settings (see Appendix D for more details of the threat model). For KD, the predicted distributions are used specifically under grey-box settings.
Metrics. For text generation tasks, we evaluate extracted models with a semantic-level and two lexical-level metrics, BERTScore (Zhang et al., 2020), BLEU (Papineni et al., 2002), and Rouge-L (Lin, 2004), all of which are commonly used in the NLG evaluation. Regarding reasoning tasks (e.g., QA), we use Precision, Recall, Accuracy, and F1 score as their evaluation metrics.
Implementation Details. We use Llama3-8B as the local model to learn the outputs generated by victim models. We set sequence length varying 128 to 4096 depending on the selected tasks, and learning rate $3\times 10^{-5}$. Our experiments run on 2 $\times$ 80GB Nvidia Tesla A100. We execute each training five times and record the mean values and standard variances. For LoRD, we set $\tau_{1}$ and $\tau_{2}$ to 0.8 and -0.1, respectively. Besides, we set the period number $N_{t}$ to 512, and use $\lambda_{1}=0.5$.
@@
Table 1: MEA comparison on WMT16 (Bojar et al., 2016) among MLE and our LoRD methods, where we use GPT-3.5-turbo as the victim model, and Llama3-8B (Grattafiori et al., 2024) as the local initial model.
@@ 5.2 Stealing Domain-Specific Knowledge
We first select GPT-3.5-turbo, a checkpoint of ChatGPT, as the basic victim model. This is because its API provides probabilities of candidate words when generating responses. We employ Llama3-8B (Grattafiori et al., 2024), a small LLM with only a 4.5% fraction of parameters than the victim model as our initial local model. Though this LaViSH (Large-Victim-Small-Heist) setting contradicts previous assumptions (Tramèr et al., 2016; Papernot et al., 2017; Jagielski et al., 2020) in MEA that the copy model should usually be “wider” or “larger” than the victim model to contain its knowledge, we believe this setting is more applicable in real world scenarios (Li et al., 2023b). Appendix D provides more detail for this setting. Besides, the number of query times selected in this section is less than 100, a significant degradation compared to previous studies (Li et al., 2023b). This is because, in our experiments, copy models can easily learn the knowledge with a few training samples and then exhibit only slight improvements afterward. More discussions on query times can be found in Appendix A.2.1.
    Model/Metric Accuracy Precision Recall F1 Score
PIQA (Bisk et al., 2020) with 64 query samples
Victim Model 0.828 0.828 0.827 0.827
Local Model 0.622 0.638 0.621 0.609
+LoRD (ours) $0.785\pm 0.01$ $0.795\pm 0.01$ $0.785\pm 0.01$ $0.783\pm 0.02$
TruthfulQA (Lin et al., 2021) with 64 query samples
Victim Model 0.414 0.500 0.207 0.293
Local Model 0.391 0.500 0.195 0.281
+LoRD (ours) $0.408\pm 0.05$ $0.500\pm 0.00$ $0.204\pm 0.03$ $0.289\pm 0.03$
Table 2: MEA comparison on QA tasks among MLE and our LoRD methods. More experiments are shown in Table 7.
@@
Fidelity and limits on stealing. We first examine the fidelity and limits of a small LLM to steal commercial LLMs. As shown in Table 1, 2 and 7, we list the performance of the victim model and the local model on five tasks, and provide two MEA methods, local model fine-tuned with MLE (+MLE) and LoRD (+LoRD), respectively.
@@
We can see that the original performance of the local model is significantly lower than the victim model, i.e., with a 50% decrease in BLEU-4 or $10\sim 25$ decrease in Rouge-L. Once we employ MEAs in the local model, its performance rapidly boosts to nearly the same as the victim model, with $0\sim 40\%$ points of gaps in BERTScore. These gaps are negligible (e.g. $<1\%$ in summarization) in some tasks, but remain eminent in other tasks such as reasoning, structured text generation, and machine translation. This phenomenon indicates that domain-specific model extractions can effectively learn domain-specific abilities from victim models but may perform poorly if downstream tasks require extra knowledge, such as machine translation and QA. We provide a stealing comparison among different local models in Table 9.
@@
Comparison among stealing methods. Tables 1, 2, and 7 compare the stealing efficacy between MLE and our LoRD. The results consistently show that LoRD outperforms MLE under the same MEA settings. Besides, for challenging tasks such as reasoning and translation, LoRD exhibits much higher improvements, which demonstrates that it can address the preference overfitting problem discussed in Section 4.2 and do enable the local model to learn the task ability from victim models. However, we also observe that for some tasks (e.g., summarization), LoRD shows no statistical difference from MLE, probably because these tasks are relatively simple, where merely MLE has already achieved comparable results to victim models.
  DiaSafety SafeRLHF
Model Toxicity Insult Profanity Severe Toxity Threat Toxicity Insult Profanity Severe Toxity Threat
Llama3-8B (initial) 14..20 7.94 8.35 1.58 2.29 7.92 2.71 2.80 0.30 1.49
+MLE 8.31 3.69 4.31 0.83 1.50 4.87 1.98 1.66 0.16 1.02
+LoRD 6.45 2.81 3.56 0.71 1.34 3.55 1.15 2.84 0.38 0.79
Table 3: Comparison on safety alignment extraction tasks.
@@
where $\mathbf{y}_{N_{t}}\sim P_{\theta_{N_{t}}}(\cdot|\mathbf{x})$, $\mathbf{y}_{0}\sim P_{\theta_{0}}(\cdot|\mathbf{x})$, and $\mathbf{y}_{vic}\sim P_{\theta_{vic}}(\cdot|\mathbf{x})$ denote the sampled responses from the trained local model ($\theta_{N_{t}}$), the initial local model ($\theta_{0}$), and the victim model ($\theta_{vic}$), respectively. In Figure 5, we illustrate a “spectrum” of extracting various downstream tasks based on these two metrics defined in Equation 12. The figure can assist in recognizing and defending commercial LLM’s knowledge. From Figure 5, we observe five tasks forming the following three scenario groups and datasets coming from the same tasks are mostly in the same group:
@@
$\bullet$ High fidelity & performance-up (HFHP). These tasks are challenging for a pre-trained model but can be effectively learned with the guidance of victim models. This group includes two tasks: data-to-text and structured text generation.
@@
$\bullet$ High fidelity & low performance-up (HFLP). The initial local model already achieves a comparable performance to the victim model. QAs and summarization are in this group.
[…]
Figure 5: Spectrum of the fidelity and performance-up on extracting different downstream tasks.
5.3 Stealing Safety Alignments
Besides of the domain-specific model extraction, we also propose the safety alignment extraction. Specifically, we select two popular safety alignment datasets for the experiments, namely SafeRLHF (Ji et al., 2024) and DiaSafety (Sun et al., 2022), to assess the safety of the generated responses. We employed PerspectiveAPI  [^2] https://perspectiveapi.com/ to automatically evaluate the safety of the responses. We select five key aspects of safety probabilities: Toxicity, Insult, Profanity, Severe Toxicity, and Threat. In these categories, a lower score indicates better safety performance. For the LoRD model, we have retained the same hyper-parameters as those used in our domain-specific experiments to ensure consistency. As shown in Table 3, we can see that both MLE and LoRD significantly reduce the harmful information after the stealing procedure. However, LoRD consistantly outperforms MLE on most of the indicators, suggesting that it can achieve better performance in the alignment task.
6 Conclusion
In this paper, we have focused on the extraction problem of commercial large language models. We proposed LoRD, a practical and realistic extraction algorithm which is consistent with the alignment procedure of large language models. Our analysis proved that LoRD can reduce the query time significantly and mitigate the certification of current watermarks naturally, surpassing existing MEA algorithms’ capabilities. Extensive experiments on domain-specific stealing and alignments demonstrated the superiority of our method.
Limitations and Future Works
MEAs on Multi-modal Models. While this paper delves into MEAs for large language models, it acknowledges the oversight of the multi-modal attribution of current commercial models (Anil et al., 2024; Achiam et al., 2024) that integrate various forms of data such as text, images, voice, and so on. The challenge of extending MEA algorithms to accommodate these models, which requires extra considerations on the unified representation of concepts, remains unexplored. Future work could focus on developing MEA methodologies sensitive to multi-modal data nuances.
Capacities beyond LaViSH Settings. We utilize the LaViSH setting to describe the model capacity of adversaries in our threat model (see Appendix D). However, sometimes, the adversary might possess comparable or superior training resources to the victims. Though this paper posits that our MEA algorithms and theoretical analysis are still compatible with such conditions, we concede that concrete experimental validation and results beyond LaViSH settings are not presented here.
Lower-level Extractions. This study evaluates MEAs at the performance level, i.e., it measures the extraction effectiveness simply through task performance metrics, or the similarity of learned distributions to the victim model. This setting is justified, as performance metrics are essential for evaluating task-related knowledge and the practical application of LLMs. However, it does not consider the lower-level similarities between the victim and local models. Can we achieve neuron-level alignments in LLM’s MEAs? How does a LaViSH setting hurt LLM’s MEAs? Is it compatible to extract a MoE (Mix-of-the-Expert) (Shazeer et al., 2017) victim model with a dense local model? These questions are not addressed in this research.
Ethical Considerations
As discussed in Section 1, MEAs are becoming increasingly prevalent in industrial settings and have already been executed, yet there remains a critical gap in understanding which specific tasks are more susceptible and what capabilities are necessary for effective executions. This lack of knowledge exacerbates the challenges faced by LLM maintainers in safeguarding their systems. Our research can contribute to that. Besides, the theoretical problem we address (as shown in Section 4) offers a novel and insightful perspective on the nature of this threat. Based on these two points, we believe the benefits of our paper outweigh potential harms, which aligns with the principles of the Menlo Report (Bailey et al., 2012) on ethics.
Additionally, we have submitted an anonymous version of the paper to the maintainers of the victim models used in our study to assist in improving their model security.
It is important to acknowledge, however, that the algorithms we propose could inadvertently enhance the efficiency of illicit extraction efforts by adversaries. To mitigate this risk, we have introduced and analyzed two defensive strategies, assessing both their effectiveness and potential vulnerabilities under adaptive attack scenarios.
Potential defenses consist of:
$\bullet$ Query Detection. One approach to effectively prevent the attack of LoRD is by detecting the distribution of query texts. This is because LoRD, similar to current MEA algorithms, makes no improvements to query samples, indicating that it can be detected by analyzing the statistical information of the adversary’s queries, such as the number of queries, distribution of query contents, and so on. However, this defense is usually resource-consuming, as it requires the LLM provider to store all query texts of each user. Besides, the potential for false positives could adversely affect the user experience.
$\bullet$ More Powerful Watermarks. While we highlight the watermark resistance of LoRD, watermarking remains one of the most effective solutions to mitigate MEAs. For example, some model-level watermarks, such as backdoor-based watermarking (Jia et al., 2021; Lv et al., 2024), can effectively certify the theft of DNNs. While model-level (e.g. backdoor-based) watermarks on pre-trained models raised increasing concerns recently (Peng et al., 2023; Gu et al., 2022; Li et al., 2023a), model-level watermarking on LLMs remains preliminary. Besides, this technique might not work when the adversary only steals a subset of knowledge in which no backdoor is embedded.
Appendix A Supplemental Experiments
A.1 Resistance to Watermarks
Current LLM watermarking methods have been shown (Kirchenbauer et al., 2023) to be robust against commonly used erasing strategies (e.g., rephrasing), making watermark removal a distinct challenge. In this section, we validate the inherent resistance of LoRD to watermarks, suggesting that LoRD is preliminarily resistant to text watermarking. As described in Section 4, we highlight that LoRD can extract the victim models’ knowledge with two terms: the straightforward likelihood learning term $\text{log}P_{\theta_{t}}(\mathbf{y}_{vic}|\mathbf{x})-\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})$ and the exploration term $\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|\mathbf{x})-\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})$, where we can tune $\lambda_{1}$ as shown in $\mathcal{L}$ to trade off the exploration and the convergence speed. Typically, a lower $\lambda_{1}$ encourages the model for conducting a slower but more diverse and localized exploration from its own generated text $\mathbf{y}_{t-1}^{+}$, potentially enhancing watermark resistance. In this subsection, we evaluate this analysis empirically.
Watermarking Details. Unlike previous experimental settings in Section 5, here we cannot utilize commercial LLMs as victim models due to the inability to control token sampling inside LLMs. Instead, we employ Llama3-70B as the victim model and watermark its outputs based on “green” tokens selection. Following prior research (Kirchenbauer et al., 2023), we separate the predicted vocabulary into a green word set and a red word set , assigning them randomly with the seed derived from the hash values of generated tokens at the last generation step. Subsequently, we sample the next token exclusively from the green set, determined by a certain probability.
Figure 6: Comparison of watermarks resistance.
@@ A.1 Resistance to Watermarks
A.2 Scaling the Stealing
In this subsection, we explore essential capacities to steal domain-specific knowledge from LLMs. We first analyze the influence of query times for the adversary, then compare the efficacy when utilizing different sizes of the local model, and finally compare the fidelity among different victim and local models.
A.2.1 Query Times
@@
A.2.2 Scales of Local Models
As shown in our threat model (see Appendix D), we assume the adversary is stealing existing commercial LLMs with a small local model. This raises the question of selecting an appropriate interval of the local model’s size. To address this concern, we illustrate the correlation between the local model’s size and extraction efficacy on two machine translation tasks, Russian-to-English (ru-en) and German-to-English (de-en), as shown in Figure 8. Here, we employ seven OPT models (Zhang et al., 2022) as local models, with parameters ranging from 125 million to 30 billion, to minimize the interruptions of factors other than model size.
@@ A.2.2 Scales of Local Models
A.2.3 Fidelity under Different Victim and Local Models